Journalist Report – February 7th
Crew 274 Journalist Report 2-7-2023
Author: Tony DiBernardo, Crew Journalist
Crew 274 woke up this morning at 7:30am to the song “There’s a Great Big Beautiful Tomorrow” from the Carousel of Progress at Walt Disney World. The crew got their surveys done, then prepared for a double EVA day utilizing 6 of the 8 crew members.
Today, the crew conducted two EVAs. The purpose of EVA #8 was to deliver a resupply of 2L of water and 6 gel packs to the supply cache, EEVAC-1. HSO Pender collected performance metrics collection and a status check on the condition of the cache and collected data on the orange radiation experiment led by Commander Sarah Guthrie. EEVAC-1 was found in good condition, however, the solar panels charging the cache had an accumulation of dried mud on them. Additionally, most of the legs supporting the solar panels had shifted, and one of them completely collapsed. According to historical weather data, it appears that winds up to 26 mph occurred in the Hanksville area the night of Sol 7. It is suspected that this weather event contributed to the state we found the solar panels in. We were able to clear the mud accumulation with gentle hand swipes and reposition the solar panel legs back to their original configuration. It is noteworthy that the stakes and ropes securing the solar panels all held up over this period, and we believe the stakes installed at the base of the solar panels played a significant role in preventing more drastic impacts.
Additionally, the crew completed the first in a series of extravehicular activities (EVAs) focused on evaluating contingency methodologies and assistive devices, or rather, how we might rescue an incapacitated astronaut on the Moon or Mars while wearing spacesuits. This study uses a Kinetic Utilization Research Tool (“K.U.R.T.”), a weighted grappling dummy; a medical evacuation sled, and an engineered vest with multiple hoist points.
The concept is based on current operation considerations for Artemis missions to the Moon requiring proper suit interface and methods for contingency EVAs. The principal investigator, Commander Guthrie, has performed this study at other analogs, but through that experience, has modified and improved the concept through lessons learned. The unique topography of the Utah desert here at MDRS provides this study with the most analogous environment to perform these tests, which could not be achieved at other analog locations.The crew looks forward to more opportunities in the coming week to rescue “KURT” on the Martian terrain.
Hab Specialist O’Hara participated in EVA 9 this afternoon where he simulated methods for recovering an incapacitated crew member in the field. After the EVA, Bill was able to complete a review of the lower level of the Habitat module. He is on track to complete his study before the conclusion of the mission.
Today, crew astronomers Pena and Loy observed detailed heliophysics phenomena. Multiple observations, videos, and photos were taken of solar dark spots, solar chromosphere granules, and solar prominences. Over 13 large solar prominences were recorded today, the largest dataset of prominences since experimentations began. These observations will be very useful for the astronomer’s analysis of solar cycle 25, how space weather implicates orbital assets, space operations planning to safeguard those assets, and the importance of having heliophysic astronomers on site at Martian Habitats.
Cloudy weather conditions and robotic observatory mechanical failures blocked the astronomy crew from from observing HADS Variable Star V0799 AUR again. Crew Astronomers instead continued their analysis of their current image inventory, stacking and analyzing 9 separate 45-second images in the visual filter. So far, HADS Variable Star V0799 AUR is showing signs of dimming, which provides good data to the astronomers’ variable star report.
No significant progress was seen from any seeds or microgreens today. GreenHab Officer Hines watered the plants in the GreenHab and continued his daily responsibilities.
EVA Report – February 7th
Crew 274 EVA Report 07-02-2023
EVA # 8
Author: Nicholas S. Pender, HSO
Purpose of EVA: Resupply, data collection, and status check of EEVAC-1 (supply cache).
Start time: 10:00 AM
End time: 11:30 AM
Narrative: EVA #8’s purpose was to deliver a resupply of 2L of water and 6 gel packs to EEVAC-1 (the supply cache) located at coordinates N4248400, E519300. Performance metrics collection and a status check on the condition of the cache was also conducted. Finally, we collected data on the orange radiation experiment led by Commander Sarah Guthrie. EEVAC-1 was found in good condition, however, the solar panels charging the cache had an accumulation of dried mud on them. Additionally, most of the legs supporting the solar panels had shifted, and one of them completely collapsed. According to historical weather data, it appears that winds up to 26 mph occurred in the Hanksville area the night of 25 Feb 23 (Sol 7). It is suspected that this weather event contributed to the state we found the solar panels in. We were able to clear the mud accumulation with gentle hand swipes and reposition the solar panel legs back to their original configuration. It is noteworthy that the stakes and ropes securing the solar panels all held up over this period, and I believe the stakes installed at the base of the solar panels played a significant role in preventing more drastic impacts. Listed below are the collected metrics:
- Inside Temp (Thermostat): 51.5 F
- Inside Temp (Data Logger): 60.6 F
- Outside Temp: 64.4 F (likely due to position in sunlight)
- Battery Level: 4 + 1 (charging)
- Power Consumption: 0W
- Total Consumption: 0.00 kWh
- Inside Temp: 61-57 F
- Outside Temp: 61 F (likely due to position in sunlight)
- Inside radiation Lvl: 0 mSv
- Outside radiation Lvl: 0 mSv
Destination: South on Cow Dung Road.
Coordinates (use UTM WGS 84): N4248400, E519300
Participants: Nicholas Pender (HSO), Lex Lojek (Engineer)
Road(s) and routes per MDRS Map: EVA #8’s (map provided) purpose is to drive to EEVAC-1 deployed at N4248400, E519300 via south on Cow Dung Road near Zubrin’s Head.
Mode of travel: Rover: Spirit
Crew 274 Sol 9 EVA Report 02-07-2023
EVA #: 9
Author: Sarah “Ceres” Guthrie, Commander
Start time: 2:00pm
Stop time: 3:30pm
Narrative: Crew performed the first of the adaptive rescue vests with medical sled EVAs at the ravine at Gateway of Candor. They dragged KURT (~80 lbs) to the destination which was physically strenuous but the sled performed well over the various terrain features (soft riverbed and rock). The crew also simulated KURT, an incapacitated who fell from the hill by placing him on the side of the hill slope. They tested the rescue vest, simulating a pressurized suit stance, lifting KURT and different points. Each member of the crew performed a single-person and two-member rescue with lifting KURT at different connection points and placing KURT on the sled. They then practiced single-person and two-member carry of the sled and walked back to the rover. The location of this test, while seeming ideal on paper, proved to be difficult in performance and further than expected. The requested time for this EVA was not sufficient to reach the location and perform the tests on the crew members. The EVA team, to our surprise, was able to maintain constant communication with the IVA (hab crew) at the EVA site. NOTE: Future EVA requests will inquire about performing more tests in this location in order to maintain comms north of the habitat. Unfortunately, the cache EVA crews seem to lose comms when they are south of the habitat.
Purpose of EVA: The purpose of this EVAs is to test the adaptive rescue vest on KURT (test dummy) with the rescue sled hilled and rocky surfaces. This location provides the scenario where incapacitated astronaut would have fallen into a ravine that can not be accessed by a rover. The crew will save “KURT” from the river bed at the base of the hill and return him to the rover parked on Cow Dung Road (N4251500, E518800).
Destination: Gateway to Candor
Coordinates (use UTM WSG 84): N4251500, E519500
EVA Participants: Sarah Guthrie (Crew Commander), Noah Loy (Heliophysics), Bill O’Hara (Crew Engineer), Tony DiBernardo (Journalist)
Road(s) and routes per MDRS Map: EVA #9 will Cow Dung Road north to Gateway to Candor.
Mode of travel: Walking and Rover (Opportunity/Curiosity)
Vehicles you will be using (If applicable): Percy & Opportunity
Astronomy Report – February 7th
Name: Salina Peña and Noah Loy
MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested (Choose one) MDRS-14 MDRS-WF
Objects Viewed: None
Problems Encountered: None. The dome MLC RCOS 16 was closed due to bad weather.
Solar Features Observed: Sun/230207/solar prominence
Problems Encountered: None (We are still processing images from today. We will send you images later.)
GreenHab Report – February 7th
GreenHab Officer: Tyler Hines
Environmental control: heater
Average temperatures: 78.15 F
Hours of supplemental light: 4 hours
Daily water usage for crops: 13.74 gallons
Daily water usage for research and/or other purposes: N/A
Water in Blue Tank: 73.63 gallons
Time(s) of watering for crops: 10:27 am
Changes to crops: N/A
Narrative: Nominally performed general watering and maintenance of plants and crops as part of daily GreenHab operations. Another notable harvest was also conducted to support upcoming crew meal planning. Further cleaning and trimming of crops were also performed as a direct result.
Harvest: 99 grams of tomato, 26 grams of lettuce, 30 grams of carrot, 26 grams of spinach, and 8 grams of kale microgreens.
Support/supplies needed: N/A
Operations Report – February 7th
Crew 274 Operations Report 2/7/2023
Name of person filing report: Alexis Lojek
Non-nominal systems: Toilet. Robotic Observatory.
Notes on non-nominal systems: Mission Support will install a new toilet holding tank between crews. Robotic observatory is closed until further notice. Please see more detailed explanation under “summary of any observatory issues.”
Spirit rover used: Yes
Beginning charge: 100%
Ending charge: 49%
Currently charging: Yes
Opportunity rover used: Yes
Beginning charge: 100%
Ending charge: 72%
Currently charging: No
Curiosity rover used: Yes
Beginning charge: 100%
Ending charge: 76%
Currently charging: No
Perseverance rover used: No
Beginning charge: 100%
Ending charge: N/A
Currently charging: Yes
General notes on rovers: Rover Spirit was used for the morning EVA and plugged in before noon to charge as it was a beautiful, cloudless day outside. Opportunity and Curiosity were used for the afternoon EVA, but returned after noon, so were not plugged in to charge.
Summary of Hab operations: Hab is in a nominal condition.
WATER USE: Mission Support facilitated a water delivery before we were able to measure our water usage.
Water (static tank): 314 gallons remaining
Static tank pipe heater (On or off): On
Static tank heater (On or off): On
Toilet tank emptied: No
Summary of internet: Nominal
Summary of suits and radios: Suits are in nominal condition and charging. Radios were charged yesterday and are off chargers. All conditions normal.
Summary of GreenHab operations: GHO Hines watered plants from 10:27 – 11:01 pm and harvested a number of vegetables.
WATER USE: 13.74 gallons used. By the GHO’s calculations, approximately 22 gallons will be remaining at the end of the mission, after Saturday morning’s watering.
Supplemental light: 4 hours (10pm-2am)
Harvest: Yes. 99 grams of tomatoes, 26 grams of lettuce, 30 grams of carrots, 26 grams of spinach, and 8 grams of kale microgreens. Total: 189 grams.
Summary of ScienceDome operations: GHO Officer Tyler Hines’ experiment is continuing to be monitored; cress microgreens have been harvested. Forty Long Duration Exposure Facility seeds were relocated to the Greenhab and are sitting on a shelf in a wet paper towel in order to further support the germination process.
Dual split: On a timer from 10pm to 7am and under Mission Support supervision.
Summary of RAM operations: No operations were conducted in the RAM.
Summary of any observatory issues: One side of the dome is not opening because the straps that provide tension for the doors were wrapped around the motor. The stress on the straps is severe enough that there is no way to release that tension to try and rewrap it. The crew has been informed that the MDRS Observatory is inoperable until further notice.
Summary of health and safety issues: N/A
Questions, concerns, and requests to Mission Support: N/A
Sol Summary Report – February 7th
Crew 274 Sol Summary Report 02-07-2023
Summary Title: Sunspots, Rescues, and Focused Breathing
Author’s name: Sarah E. Guthrie (“Ceres”), Commander
Mission Status: Nominal
Sol Activity Summary: The crew performed the first in a series of rigorous EVAs which evaluate the use of rescue of a vest which hopes to assist lifting an incapacitated astronaut from the surface. The vest was engineered with multiple lift points for the ease of a one person rescue or two-member team. Additionally, the test uses a medical sled to transport the incapacitated astronaut of various terrain. The astronomy team captured amazing solar images with dozens of sunspots and prominences, unfortunately mechanical failure has closed the remote telescope for night time observations at MDRS leaving the team to use a different remote telescope located in New Mexico. As noted, this crew is particularly larger with 8 members and Mission Support kindly supported the team with an additional replenishment of water. Starting on Sol 7, the crew will continue to meditate in the evenings as part of Engineer Lojek’s study which hopes to reduce stress in high stress moments.
Look Ahead Plan: The crew looks forward two completing more rescue EVAs in the final days of this mission(pending approval).
Anomalies in work: None
Weather: Sunny, 1C High, -3C Low; Sunny
Crew Physical Status: Energized
EVA: EVA #9 completed for incapacitation rescue – successful
Reports to be filed: Sol Summary, Operations, EVA Request, EVA Report, GHO Report, and Astronomy Report.
Support Requested: None
You must be logged in to post a comment.